OpenStudy (erniellyd28):

x + y ------ z when x = 2 y = 3 z = 1 --- --- ---- 3 4 12

OpenStudy (baldymcgee6):

(x+y)/z, when x =2/3, y=3/4, z=1/12 How would you approach this problem?

OpenStudy (erniellyd28):

do u add first?

OpenStudy (baldymcgee6):